# Is Causality Necessary for Efficient Portfolios? A Computational Perspective on Predictive Validity and Model Misspecification

### v3

@@ 3.1.3 Conditions Underpinning the Necessity Claim
The central thesis of (López de Prado et al.,, 2025) is that causal factor modeling is a necessary condition for investment efficiency. However, this claim rests on several assumptions, some explicit, others implicit, which are listed and critically evaluated in Table 1.
@@
C1 Causal factor models are necessary for portfolio efficiency. False — disproven by counterexample.
@@
C5 All observed associations must be interpreted causally; correlation cannot justify predictive modeling. False — contradicted by predictive theory (see (Shmueli,, 2010)).
@@
Table 1: Conditions required for the necessity of causal modeling, as claimed in (López de Prado et al.,, 2025), and prior work (López de Prado et al.,, 2024; López de Prado and Zoonekynd,, 2024).
@@ 3.2 Structural Cancellation as a Counterexample
We begin with a counterexample showing that omitted variables do not inevitably cause signal inversion. Instead, omitted-variable bias can produce approximate cancellation, preserving directional validity while attenuating magnitudes. This challenges prior claims that structural misspecification necessarily yields inverted signals and collapsed frontiers.
@@ 3.3 Optimization Under Misspecification
We now distinguish prediction, ranking, and efficient allocation. Our results demonstrate that directional agreement between signals is necessary but not sufficient for portfolio efficiency: signals may align in sign yet still yield inefficient allocations due to miscalibration. This refines prior claims that inefficiency arises only through sign inversion.
@@
$R(\omega_{\mathrm{pred}})\neq R(\omega_{\mathrm{true}}),\quad\mathrm{Sharpe}(\omega_{\mathrm{pred}})<\mathrm{Sharpe}(\omega_{\mathrm{true}}),$
where $R(\omega)=\mathbb{E}[\omega(x)r(x)]$ and $\mathrm{Sharpe}(\omega)=\tfrac{\mathbb{E}[\omega(x)r(x)]}{\sqrt{\mathrm{Var}(\omega(x)r(x))}}$.
@@
$R(\omega_{\mathrm{pred}})=\epsilon R(\omega_{\mathrm{true}}),\quad\mathrm{Sharpe}(\omega_{\mathrm{pred}})=\epsilon\,\mathrm{Sharpe}(\omega_{\mathrm{true}}).$
Directionality is preserved, but performance is strictly attenuated when $\epsilon\neq 1$. The case $\epsilon<0$ recovers full inversion, a special case of inefficiency previously emphasized. ∎
Remark. Directional correctness ensures that signs match, but without proper scaling, efficiency contracts.
@@
Corollary 2 (Calibration Is Necessary for Efficiency).
@@
Let $\mu$ be the true mean return vector, $\tilde{\mu}$ a surrogate preserving its order, and $V\succ 0$ the covariance matrix. Then
@@
where $\rho_{V^{-1}}=\tfrac{\mu^{\top}V^{-1}\tilde{\mu}}{\|\mu\|_{V^{-1}}\|\tilde{\mu}\|_{V^{-1}}}\in[-1,1]$.
@@
By Almgren and Chriss, (2005), order preservation ensures $V^{-1}\tilde{\mu}$ lies in the efficient cone. Computing return and variance under $V^{-1}$ geometry yields the stated scaling law. ∎
@@ 3.4 Testing Optimization Geometry Under Misspecification
We now examine whether structural misspecification necessarily collapses portfolio optimization geometry. Using nonlinear confounded processes, we show that surrogate signals, even when estimated from misspecified models, can yield viable efficient frontiers provided they remain directionally aligned with the true return vector.
@@ 3.4.1 Nonlinear Confounded Signals
$Z=\alpha_{1}X_{1}+\alpha_{2}X_{2}+\eta,\quad\eta\sim\mathcal{N}(0,1).$
@@
$\mathbb{P}(Y=1\mid X,Z)=\sigma\!\big{(}\tanh(X)+0.5\sin(Z)\big{)},$
@@
$\omega^{\mathrm{true}}=2\,\mathbb{P}(Y=1\mid X,Z)-1.$
@@
$\omega^{\mathrm{pred}}=2\,\sigma(\hat{\beta}^{\top}X)-1.$
@@ 3.4.2 Robustness of Optimization Geometry
$\max_{\omega\in\mathbb{R}^{n}}\quad\tilde{\mu}^{\top}\omega-\lambda\,\omega^{\top}\Sigma\omega\quad\text{s.t. }\mathbf{1}^{\top}\omega=1,\ \lambda>0.$
@@
$\omega^{*}(\lambda)=\tfrac{1}{2\lambda}\Sigma^{-1}(\tilde{\mu}-\nu\mathbf{1}),$
@@
Immediate from Lemma 4, with $\tilde{\mu}$ interpreted as a surrogate mean vector. Directional alignment ensures positive realized returns, convexity, and smooth variation of $\omega^{*}(\lambda)$. ∎
Remark. Structural misspecification does not collapse optimization geometry. Directional informativeness suffices to preserve frontier viability.
@@ 3.4.3 Sharpe Ratio Sensitivity to Alignment
Let $\mu\in\mathbb{R}^{n}$ denote true returns and $\nu\perp\mu$ with $\|\nu\|=\|\mu\|$. Define
$\tilde{\mu}(\theta)=\cos\theta\,\mu+\sin\theta\,\nu,\quad\theta\in[0,\pi].$
For $\omega^{*}(\theta)=\tfrac{1}{2\lambda}\Sigma^{-1}\tilde{\mu}(\theta)$, the realized Sharpe ratio satisfies
@@
$\mu^{\top}\omega^{*}(\theta)=\tfrac{1}{2\lambda}\mu^{\top}\Sigma^{-1}\tilde{\mu}(\theta)=\cos\theta\cdot\mu^{\top}\omega^{*}(0),$
since $\mu^{\top}\Sigma^{-1}\nu=0$. Portfolio volatility $\sigma(\omega^{*}(\theta))=\tfrac{1}{2\lambda}\sqrt{\tilde{\mu}(\theta)^{\top}\Sigma^{-1}\tilde{\mu}(\theta)}$ is invariant in $\theta$ by construction. Dividing return by volatility yields the stated scaling. ∎
@@
  • $\cos\theta=0$: the portfolio is orthogonal to $\mu$ and achieves zero return.
@@ 3.5 Approximate Cancellation and Attenuation
We study omitted‐variable bias when a confounder nearly cancels the observed signal but is corrupted by independent noise. Let $X\sim\mathcal{N}(0,1)$, and let
@@ 4 Numerical Results
We now provide empirical validation of the theoretical framework developed in Section 3, focusing on the geometry of portfolio optimization under structural misspecification. The experiments are organized in five parts. First, simulations illustrate structural cancellation, showing that omitted-variable bias may attenuate but not invert signals (Corollary 1). Second, calibration–ranking experiments validate that ranking ensures efficient-set membership, whereas calibration governs quantitative efficiency (Lemma 3, Proposition 1). Third, nonlinear counterexamples test robustness of optimization geometry, demonstrating that misspecification does not collapse the frontier (Lemma 4, Corollary 3). Fourth, alignment experiments confirm that Sharpe ratios contract linearly with cosine alignment, as predicted by Proposition 2 and Corollary 4. Finally, empirical validation with S&P 500 equity data shows that predictive signals from misspecified models can sustain convex and coherent frontiers in practice.
Together, these results disentangle the distinct roles of directionality, ranking, and calibration in determining portfolio outcomes. All simulation code is available in the supplementary material.
@@ 4.1 Structural Cancellation Under Misspecification
Figure 1: Portfolio Weights under Structural Cancellation. Scatter of true weights (x-axis) against misspecified weights (y-axis). Signs are preserved but magnitudes are attenuated, showing directional consistency despite misspecification.
@@ 4.2 Calibration and Ranking within the Efficient Set
(a) Ranking preserved under monotone transforms ($p=0.6,0.8,1.0$). Order is maintained across assets, ensuring efficient-set membership.
(b) Sharpe ratios as a function of $p$. Efficiency peaks at $p=1$ and declines smoothly for $p\neq 1$, consistent with Proposition 1.
Figure 2: Calibration and Ranking Experiments. Monte Carlo validation of Section $3.$ Panel (a): monotone transforms preserve ranking, consistent with Lemma 3. Panel (b): Sharpe efficiency contracts with miscalibration, confirming Proposition 1. Together, these results show that ranking ensures validity, while calibration determines frontier position.
@@ 4.3 Nonlinear Counterexamples and Frontier Geometry
Figure 3: Misspecified vs. True Portfolio Weights. Scatter of predicted versus true weights under a nonlinear confounded model. Misspecification compresses magnitudes but does not induce systematic sign inversion.
Figure 4: Efficient Frontier from Misspecified Signals. Portfolios based on misspecified signals yield a smooth, convex frontier, illustrating preserved geometry despite reduced efficiency.

### v4

@@ 3.1.3 Conditions Underpinning the Necessity Claim
The central thesis of López de Prado et al. (2025a) is that causal factor modeling is a necessary condition for investment efficiency. However, this claim rests on several assumptions, some explicit, others implicit, which are listed and critically evaluated in Table 1.
@@
C1 Causal factor models are necessary for portfolio efficiency. False within the static optimization framework studied here — disproven by counterexample.
@@
C5 All observed associations must be interpreted causally; correlation cannot justify predictive modeling. False — contradicted by predictive theory (see (Shmueli, 2010)).
@@
Table 1: Conditions required for the necessity of causal modeling, as claimed in López de Prado et al. (2025a), and prior work López de Prado et al. (2024); López de Prado and Zoonekynd (2024).
@@ 3.2 Structural Cancellation as a Counterexample
This section addresses the claim that structural misspecification generically leads to signal inversion. We begin with a counterexample showing that omitted variables do not inevitably cause signal inversion. Instead, omitted-variable bias can produce approximate cancellation, preserving directional validity while attenuating magnitudes. This challenges prior claims that structural misspecification necessarily yields inverted signals and collapsed frontiers.
This counterexample establishes structural robustness to misspecification: omitted-variable bias may attenuate predictive signals without inducing systematic sign inversion, thereby preserving directional informativeness under moderate departures from the true data-generating process.
@@ 3.2.2 Misspecified Estimation
In econometric analysis, omitted-variable bias is typically assessed through consistency and causal interpretability, where bias implies invalid inference. In portfolio optimization, however, the relevant criterion is directional informativeness for allocation, rather than unbiased estimation for inference. The results above show that omitted-variable bias may attenuate or rescale predictive signals without destroying their alignment with realized returns, explaining why structural misspecification can invalidate causal interpretation while leaving optimization performance intact. This distinction parallels the explanatory–predictive divide emphasized by Shmueli (2010) and Breiman (2001), where models may be unsuitable for causal interpretation yet remain operationally effective for prediction.
@@ 3.3 Optimization Under Misspecification
This section addresses the claim that directional correctness alone is sufficient for portfolio efficiency. We now distinguish prediction, ranking, and efficient allocation. Throughout this section, efficiency initially refers to optimality with respect to the mean–variance objective, and later to position along the mean–variance efficient frontier. Our results demonstrate that directional agreement between signals is necessary but not sufficient for portfolio efficiency: signals may align in sign yet still yield inefficient allocations due to miscalibration. This refines prior claims that inefficiency arises only through sign inversion.
@@
$R(\omega_{\mathrm{pred}})\neq R(\omega_{\mathrm{true}}),\quad J(\omega_{\mathrm{pred}})<J(\omega_{\mathrm{true}}),$
where $R(\omega)=\mathbb{E}[\omega(x)r(x)]$ and
$J(\omega)=\mathbb{E}[\omega(x)r(x)]-\lambda\,\mathrm{Var}(\omega(x)r(x)),\qquad\lambda>0.$
@@
$R(\omega_{\mathrm{pred}})=\epsilon R(\omega_{\mathrm{true}})\neq R(\omega_{\mathrm{true}}).$
Moreover, letting $Y=\omega_{\mathrm{true}}(x)r(x)$,
$J(\omega_{\mathrm{pred}})=\mathbb{E}[\epsilon Y]-\lambda\,\mathrm{Var}(\epsilon Y)=\epsilon\,\mathbb{E}[Y]-\lambda\,\epsilon^{2}\,\mathrm{Var}(Y)<\mathbb{E}[Y]-\lambda\,\mathrm{Var}(Y)=J(\omega_{\mathrm{true}}),$
since $\epsilon\in(0,1)$ and $\lambda>0$. The case $\epsilon<0$ recovers full inversion, a special case previously emphasized. ∎
Remark. Directional correctness ensures that signs match, but without proper scaling, mean–variance efficiency contracts.
@@
Corollary 2 (Calibration Is Necessary for Quantitative Efficiency Within the Efficient Set).
@@
Let $\mu$ be the true mean return vector, $\tilde{\mu}$ a surrogate preserving its order, and $V\succ 0$ the covariance matrix. Portfolio weights are formed as $V^{-1}\tilde{\mu}$, while performance is evaluated under the true mean $\mu$. Then
@@
where
$\rho_{V^{-1}}=\frac{\mu^{\top}V^{-1}\tilde{\mu}}{\|\mu\|_{V^{-1}}\|\tilde{\mu}\|_{V^{-1}}}\in[-1,1].$
@@
We work under the standard admissibility conditions of mean–variance optimization: $V\succ 0$ and surrogate expected returns $\tilde{\mu}$ that preserve the componentwise ordering of the true mean return vector $\mu$. Under these conditions, Almgren and Chriss (2005) show that $V^{-1}\tilde{\mu}$ lies in the efficient cone. Computing return and variance under the $V^{-1}$ geometry yields the stated scaling law. ∎
Proposition 1 differs from standard mean–variance results in that the expected return vector is not assumed to be correctly specified. Classical treatments implicitly treat $\mu$ as a primitive object, whereas the present result characterizes portfolio efficiency when $\tilde{\mu}$ is an informational surrogate whose alignment with $\mu$ may vary continuously. This distinction is central to the causal-necessity debate: efficiency does not hinge on structural correctness, but on the geometric relation between surrogate and true signals. This attenuation effect is consistent with the well-documented sensitivity of mean–variance portfolios to expected return estimation error (Best and Grauer, 1991; Chopra and Ziemba, 1993), but differs in that it characterizes efficiency loss geometrically rather than statistically.
The linear dependence of the Sharpe ratio on cosine alignment implies performance robustness: efficiency degrades smoothly with declining alignment rather than collapsing abruptly under misspecification.
@@ 3.4 Testing Optimization Geometry Under Misspecification
This section addresses the claim that structural misspecification necessarily collapses portfolio optimization geometry, yielding degenerate or ill-defined efficient frontiers. Using nonlinear confounded data-generating processes, we examine whether misspecification generically invalidates optimization. We show that surrogate signals, even when estimated from misspecified models, can sustain smooth and non-degenerate efficient frontiers provided they remain directionally aligned with the true return vector.
The relevance of nonlinear data-generating processes in this context lies not in functional form per se, but in their effect on directional alignment. Nonlinear transformations and interactions may distort magnitudes or induce heteroskedasticity, yet they do not generically imply sign reversal. As long as predictive models preserve monotone relationships with expected returns, the geometric conditions underpinning optimization (directional alignment and ranking preservation) remain intact. This perspective is consistent with empirical evidence that predictive models can outperform structurally motivated ones in high-dimensional and nonlinear settings (DeMiguel et al., 2009; Gu et al., 2020).
@@ 3.4.1 Nonlinear Confounded Signals
$Z=\alpha_{1}X_{1}+\alpha_{2}X_{2}+\eta,\qquad\eta\sim\mathcal{N}(0,1).$
@@
$\mathbb{P}(Y=1\mid X,Z)=\sigma\!\left(\tanh(X)+0.5\sin(Z)\right),$
@@
$\omega_{\mathrm{true}}=2\,\mathbb{P}(Y=1\mid X,Z)-1.$
@@
$\omega_{\mathrm{pred}}=2\,\sigma(\hat{\beta}^{\top}X)-1.$
@@ 3.4.2 Robustness of Optimization Geometry
$\max_{\omega\in\mathbb{R}^{n}}\;\tilde{\mu}^{\top}\omega-\lambda\,\omega^{\top}\Sigma\omega\quad\text{s.t.}\quad\mathbf{1}^{\top}\omega=1,\ \lambda>0.$
@@
$\omega^{*}(\lambda)=\frac{1}{2\lambda}\Sigma^{-1}(\tilde{\mu}-\nu\mathbf{1}),$
@@
Remark. The expression $\Sigma^{-1}\tilde{\mu}$ should be interpreted as an efficient direction. The budget constraint $\mathbf{1}^{\top}\omega=1$ is enforced by the affine normalization encoded by the Lagrange multiplier $\nu$, which ensures $\mathbf{1}^{\top}\omega^{*}(\lambda)=1$ and does not affect frontier geometry.
@@
Immediate from the previous lemma, with $\tilde{\mu}$ interpreted as a surrogate mean vector. Directional alignment ensures positive realized returns, convexity, and smooth variation of $\omega^{*}(\lambda)$. ∎
Remark. Structural misspecification does not collapse optimization geometry. Directional informativeness suffices to preserve frontier viability. In this sense, robustness refers to preservation of optimization geometry under misspecified inputs, not to invariance of optimal weights or realized returns.
@@ 3.4.3 Sharpe Ratio Sensitivity to Alignment
Let $\mu\in\mathbb{R}^{n}$ denote true expected returns and let $\nu\perp\mu$ with $\|\nu\|=\|\mu\|$. Define
$\tilde{\mu}(\theta)=\cos\theta\,\mu+\sin\theta\,\nu,\qquad\theta\in[0,\pi].$
For $\omega^{*}(\theta)=\frac{1}{2\lambda}\Sigma^{-1}\tilde{\mu}(\theta)$, the realized Sharpe ratio satisfies
@@
$\mu^{\top}\omega^{*}(\theta)=\frac{1}{2\lambda}\mu^{\top}\Sigma^{-1}\tilde{\mu}(\theta)=\cos\theta\cdot\mu^{\top}\omega^{*}(0),$
since $\mu^{\top}\Sigma^{-1}\nu=0$. Portfolio volatility
$\sigma(\omega^{*}(\theta))=\frac{1}{2\lambda}\sqrt{\tilde{\mu}(\theta)^{\top}\Sigma^{-1}\tilde{\mu}(\theta)}$
is invariant in $\theta$ by construction. Dividing return by volatility yields the stated scaling. ∎
@@
  • $\cos\theta=0$: the portfolio is orthogonal to $\mu$ and achieves zero return;
@@ 3.5 Approximate Cancellation and Attenuation
This section addresses the claim that portfolio efficiency is fragile to misspecification, in the sense that small departures from the true data-generating process necessarily induce abrupt performance collapse. We study omitted‐variable bias when a confounder nearly cancels the observed signal but is corrupted by independent noise. Let $X\sim\mathcal{N}(0,1)$, and let
@@
We emphasize that robustness to misspecification is quantified explicitly rather than asserted qualitatively. Proposition 3 and Corollaries 56 establish a continuous and bounded attenuation of predictive signals as misspecification increases, while Proposition 4 shows that the resulting portfolio performance degrades smoothly through directional misalignment. Together, these results rule out threshold effects or knife-edge behavior. As a consequence, no minimum cosine similarity is required for viability; instead, efficiency degrades smoothly under increasing misspecification.
@@ 4 Numerical Results
We now provide empirical validation of the theoretical framework developed in Section 3, focusing on the geometry of portfolio optimization under structural misspecification. The experiments are organized in five parts. First, simulations illustrate structural cancellation, showing that omitted-variable bias may attenuate but not invert predictive signals (Corollary 1). Second, calibration–ranking experiments validate that ranking ensures efficient-set membership, while calibration governs quantitative efficiency (Lemma 3, Proposition 1). Third, nonlinear counterexamples test the robustness of optimization geometry, demonstrating that misspecification does not collapse the efficient frontier (Lemma 4, Corollary 3). Fourth, alignment experiments confirm that Sharpe ratios contract linearly with cosine alignment, as predicted by Proposition 2 and Corollary 4). Finally, empirical validation using S&P 500 equity data shows that predictive signals derived from misspecified models can sustain convex and coherent frontiers in practice.
Appendix A contains detailed descriptions of the simulation protocol, data preprocessing, normalization choices, and additional robustness checks. These materials support the empirical results in this section while keeping the main exposition focused on interpretation. Additional robustness analyses, including high-dimensional stress tests, rolling-window evaluations, sensitivity checks under alternative dependence and noise regimes, and constraint-aware optimization, are reported in Appendix E. Complementing the equity-based empirical analysis presented in this section, Appendix E further reports a large-scale empirical validation using a global bond universe comprising 1,350 instruments across multiple currencies, countries, sectors, maturities along the term structure, seniority classes, and credit ratings, observed over 1,061 trading days (from Bloomberg). These experiments extend the theoretical predictions to heterogeneous risk profiles, dependence structures, and realistic portfolio constraints beyond equity markets.
Together, these results disentangle the distinct roles of directionality, ranking, and calibration in determining portfolio outcomes. All simulation code is available in the supplementary material. Algorithm 1 summarizes the simulation design used throughout this section, highlighting how alignment, ranking, and calibration of predictive signals affect mean–variance portfolio optimization under structural misspecification. The simulations in this section examine empirical robustness by assessing whether directional alignment, calibration effects, and optimization geometry persist under increasing misspecification and noise.
Algorithm 1 Simulation procedure for assessing predictive signal viability
1: Generate true expected returns $\mu\in\mathbb{R}^{N}$ and covariance matrix $\Sigma\in\mathbb{R}^{N\times N}$
2: Construct a surrogate predictive signal $\hat{\mu}$ under controlled misspecification (e.g., omitted variables or nonlinear transformations)
3: Compute directional alignment $\cos(\hat{\mu},\mu)$ and assess ranking preservation between $\hat{\mu}$ and $\mu$
[…]
5: Evaluate the resulting efficient frontier and realized performance metrics (return, volatility, Sharpe ratio)
6: Repeat across varying degrees of misspecification and dependence to assess robustness
@@ 4.1 Structural Cancellation Under Misspecification
Figure 2: Portfolio Weights under Structural Cancellation. Scatter of optimal portfolio weights obtained from the true expected return vector (x-axis) against weights obtained from a structurally misspecified signal (y-axis). Each point corresponds to an asset. The preservation of sign across weights indicates directional alignment, while attenuation of magnitudes reflects miscalibration. This validates that misspecification need not invalidate diversification, but primarily reduces achievable risk-adjusted returns.
@@ 4.2 Calibration and Ranking within the Efficient Set
(a) Ranking preservation under monotone transformations. Asset ranks induced by predictive signals after monotone transformations with parameters $p=0.6,0.8,1.0$. The x-axis reports asset rank under the true signal and the y-axis under the transformed signal. Exact rank preservation across transformations implies efficient-set membership, illustrating that feasibility depends on ordering rather than magnitude accuracy.
(b) Effect of calibration on risk-adjusted performance. Population Sharpe ratios as a function of the calibration parameter pp. The x-axis indexes the degree of signal scaling and the y-axis reports the resulting Sharpe ratio. Efficiency peaks at $p=1$ (correct calibration) and declines smoothly as $p$ departs from unity, confirming the cosine-scaling prediction of Proposition 1.
Figure 3: Calibration and ranking under structural misspecification. Monte Carlo validation of the theoretical results in Section 3.3. Panel (a) shows that monotone transformations preserve asset ranking, ensuring feasibility and efficient-set membership (Lemma 3). Panel (b) shows that miscalibration alone suffices to attenuate risk-adjusted performance, with Sharpe ratios contracting smoothly as calibration deteriorates (Proposition 1). Together, the panels demonstrate that ranking secures validity of optimization, while calibration determines the attainable position along the efficient frontier.
@@ 4.3 Nonlinear Counterexamples and Frontier Geometry
Figure 4: Directional robustness under nonlinear misspecification.Scatter of portfolio weights implied by the true return signal (x-axis) versus weights implied by a misspecified predictive signal (y-axis) under a nonlinear confounded data-generating process. Each point corresponds to an asset. While nonlinear misspecification compresses weight magnitudes, the dominant diagonal structure indicates preservation of sign and relative direction. This confirms that nonlinear confounding does not generically induce signal inversion and that directional alignment can persist even when structural correctness fails.
Figure 5: Efficient frontier viability under misspecified signals. Mean–variance efficient frontier constructed using misspecified predictive signals. The x-axis reports portfolio volatility and the y-axis expected return implied by the surrogate signal. Despite reduced slope and attainable Sharpe ratios relative to the correctly specified case, the frontier remains smooth, convex, and well defined. This illustrates that structural misspecification degrades efficiency quantitatively through calibration effects but does not collapse optimization geometry.